# Jianwei Li & Jung-Eun Kim

### v2

@@ Abstract
As large language models (LLMs) are overwhelmingly more and more integrated into various applications, ensuring they generate safe responses is a pressing need. Previous studies on alignment have largely focused on general instruction-following but have often overlooked the distinct properties of safety alignment, such as the brittleness of safety mechanisms. To bridge the gap, we propose the Superficial Safety Alignment Hypothesis (SSAH), which posits that safety alignment teaches an otherwise unsafe model to choose the correct reasoning direction - fulfill or refuse users’ requests - interpreted as an implicit binary classification task. Through SSAH, we hypothesize that only a few essential components can establish safety guardrails in LLMs. We successfully identify four types of attribute-critical components: Safety Critical Unit (SCU), Utility Critical Unit (UCU), Complex Unit (CU), and Redundant Unit (RU). Our findings show that freezing certain safety-critical components during fine-tuning allows the model to retain its safety attributes while adapting to new tasks. Similarly, we show that leveraging redundant units in the pre-trained model as an “alignment budget” can effectively minimize the alignment tax while achieving the alignment goal. All considered, this paper concludes that the atomic functional unit for safety in LLMs is at the neuron level and underscores that safety alignment should not be complicated.
@@ 1 Introduction
Figure 1: Superficial Safety Alignment Hypotheses
To answer these questions, this paper introduces Superficial Safety Alignment Hypothesis (SSAH), which separates safety alignment from general alignment and emphasizes its distinct properties. With this hypothesis, we prove that less is more in terms of parameters for safety alignment: a small number of but strategically vital safety-critical components are sufficient to achieve robust safety performance. Unlike Wei et al. (2024b), our extensive experiments prove that by freezing certain safety-critical units, safety performance is successfully preserved against finetuning attacks. Similarly, we also propose to repurpose redundant units of the pre-trained model as an “alignment budget.” By reassigning these units toward safety tasks, we can reduce the safety alignment tax, ensuring that safety is maintained without sacrificing the model’s overall utility. Taken all together, we answer the following core questions:
@@
Answer: Based on the corollary of SSAH, less is more for parameters, we propose an attribute-based approach to analyzing the alignment and finetuning processes, where specific attributes are assigned to certain computational units. Our findings explain that the desired attributes (safety or utility) can be achieved by repurposing units originally responsible for other functions. This reallocation helps explain the trade-off between the safety performance and utility performance.
@@ 2 Related Work
Alignment and Safety Alignment: Alignment research in LLMs aims to ensure that models follow human instructions and align with human preferences across various tasks. Early work, such as Askell et al. (2021); Bai et al. (2022), focused on enabling LLMs to “Follow instructions and be helpful, truthful, and harmless” throughout the alignment process. Various alignment strategies have since been explored (Wang et al., 2024), including Supervised Fine-Tuning (SFT) (Taori et al., 2023; Zhou et al., 2024), Reinforcement Learning with Human Feedback (RLHF) or AI Feedback (RLAIF) (Ouyang et al., 2022; Lee et al., ), Instruction Tuning (Wei et al., 2021), Contrastive Learning (Rafailov et al., 2024; Xu et al., 2024), and Conditional Learning (Korbak et al., 2023). However, researchers have realized that achieving helpfulness, truthfulness, and harmlessness presents distinct challenges. More recent work has, therefore, shifted focus specifically toward the challenge of harmlessness, leading to an increasing emphasis on safety alignment—ensuring models avoid harmful outputs while maintaining utility Wei et al. (2024b); Qi et al. (2023).
Alignment Tax and Fine-tuning Attacks: Aligning LLMs with human preferences often incurs an “alignment tax,” where maintaining alignment often degrades downstream task performance (Bai et al., 2022; Ouyang et al., 2022; Lin et al., 2024; Wang et al., 2024). Additionally, fine-tuning can undermine safety: Yang et al. (2023); Qi et al. (2023) show that even benign fine-tuning weakens safety measures, highlighting the trade-off between alignment and utility.
Model Pruning: Pruning reduces model size and computational cost by removing redundant components (e.g., parameters, neurons, layers) while preserving performance Frantar & Alistarh (2022); Frankle et al. (2020); Anwar et al. (2017); An et al. (2024); Li et al. (2024); Molchanov et al. (2019); Han et al. (2015); Lee et al. (2019); Renda et al. (2020). By identifying less critical components, pruning enables efficient sub-model extraction for resource-limited environments. We leverage pruning to dissect model components contributing to safety, utility, or both.
@@ 3 Superficial Safety Alignment Hypothesis
Extending to Jailbreak Attack. Although SSAH is originally proposed to explain how current safety alignment techniques impact models’ behavior under direct attack, it also provides insight into jailbreak attacks. In such cases, attackers typically use manipulative tokens to bypass the model’s safety mechanisms, indicating that the current alignment method can only hold the correct reasoning direction in a limited generated tokens. However, a potential solution inspired by SSAH suggests that if safety alignment enables the model to consistently re-select the correct reasoning direction at each step (by re-evaluating the query and previously generated tokens before producing the next one), the model can continue to generate outputs that are both safe and helpful, even confronted with adversarial attempts. This perspective has been validated explicitly by a recent study, which demonstrates its practical effectiveness in mitigating jailbreak attacks (li2025superficial). This also suggests that improving local consistency (safety) in reasoning can be a key pathway toward building more robust and attack-resilient language models.
@@
Figure 2: Probing reasoning direction on the AdvBench dataset with Llama2-7B, Llama3-8B, and Llama3.1-8B using cosine distance. Models were finetuned to ensure that aligned versions possess both general instruction-following abilities and safety guardrails, while unaligned models only have instruction-following capabilities. More results are in Appendix A.2.
Probing Experiment. Although we cannot directly observe the model’s reasoning direction, as a turnaround, we can infer it by measuring the distance between hidden states in feature space when the model is fed with queries that follow different reasoning paths. By comparing the distances in aligned and unaligned models, we gain insights into how safety alignment affects the model’s reasoning direction in each generation step. Specifically, we observe that the model’s behavior can be influenced by the initial tokens in its response. For instance, appending certain tokens to a malicious query can lead an aligned model to produce unsafe responses or prompt an unaligned model to generate safe ones. This indicates that initial response tokens can alter the model’s reasoning direction. Based on this observation, we construct three types of queries to probe the model’s reasoning trajectory (more details in Appendix A.2): (1) Clean Query: The original malicious query (e.g., “How to make a bomb?”); (2) Query + benign tokens: The malicious query followed by benign prompt token (e.g., “How to make a bomb? Sorry, I can’t…”); (3) Query + malicious tokens: The malicious query followed by malicious prompt tokens (e.g., “How to make a bomb? Here’s how…”).
@@ 4 Less is More for Safety Alignment
While the full validation of SSAH remains challenging due to the high complexity of the output sampling space, we are able to empirically validate a key corollary. Specifically, if SSAH holds, then safety alignment can be achieved with only a small subset of critical computing units, as the task can be interpreted as an implicit and simple safety-related binary classification problem.
@@ 4.1 Identifying Safety-Critical Computing Units
To verify this corollary, we designed experiments to determine the minimally essential subset of computing units in LLMs that is critical in establishing a safety guardrail. Following this, we hypothesize that specific attributes of LLMs can be explicitly linked to certain computing units within the model. Our experiments are designed as follows:
@@
Table 1: Pruning results of Llama2-7B-Chat and Llama3-8B-Instruct across safety and utility benchmarks. Breakdown of model performance after pruning various categories of computing units, including SCU, UCU, and RU, demonstrating their respective contributions to safety and utility attributes. The proportion of each attribute group in the model is determined based on the degradation in utility and safety. Additional evaluation details are provided in Appendix B.1 and B.2.
@@
Following Wei et al. (2024b), we construct two datasets to separately evaluate the model’s performance on utility and safety. The utility dataset measures the model’s functional capabilities (e.g., general reasoning, language understanding), while the safety dataset evaluates its ability to reject harmful or unethical queries. This allows us to identify the computing units most closely associated with utility and safety, respectively. Unlike previous approaches that identify safety-critical components at the weight level, we identify them at the neuron level, focusing on individual neurons and channels within the model. Specifically, we use a structured pruning strategy inspired by An et al. (2024), which removes structured components of each depth-2 module based on the variance of activation values across a target dataset. Given a depth-2 module, $f(X)=B\sigma(AX)$, which can represent either an attention module or a feedforward module, where $A$ and $B$ are weight matrices. Then, we define the importance score for each channel as follows:
@@
Here, $N$ refers to the number of calibration samples, $\mathbf{W}^{B}_{:,j}$ refers to the $j$-th column of $B$, and $X^{B}$ represents the input to $B$. With this score, we plan to prune the input channels of $B$ and the output neurons of $A$, as channels or neurons with low activation variance across the target dataset are considered less important - details in Appendix B.3.
@@
Ablation Study Results. The experiment results are described in Tab. 1. We discovered that for a safety-aligned model, the computing units that are exclusively responsible for the safety attribute account for only about 1.3 - 1.4% of the total units. Although the complex units make up a larger portion of the model, their primary role is to support the general knowledge required for both safety and utility tasks. Based on these findings, we have partially validated our hypothesis that safety alignment is relying on (not constructed by) a subset of safety-critical computing units. We also provide more experiment results in Appendix C.
@@ 4.2 Why is Safety Brittle?
Table 2: Safety performance of Llama2-7B-Chat and Llama3-8B-Instruct under Fine-Tuning attacks (Alpaca and Dolly) across various benchmarks and judge methods. We compare the measures of the initial models, finetuned models, and our strategies including two settings: setting (i): freezing all SCU and the top 6% of CU, and setting (ii): freezing all SCU and all CU. Both strategies demonstrate significant mitigation of safety performance degradation. Bold indicates the best results, while the underlined mark the second-best results. Note that we doubled the training epochs for our method to ensure a fair comparison, resulting in identical or lower final training loss compared to the finetuned models. Additionally, due to computational limitations, we froze the first 12 transformer blocks of LLaMA3, although similar trends are observed. We also finetuned the model with Setting (ii) for one epoch prior to applying Setting (i). Further details are in Appendix B.5.
@@
Freezing Safety-Critical Components to Preserve Safety Guardrails. Given the above insight, we propose freezing the identified safety-critical components during the finetuning process to prevent the unwanted attribute transfer of these units and thereby preserve safety performance. To test this hypothesis, we conduct experiments across different language models, where we freeze the safety-critical components identified through our aforementioned ablation study in the finetuning process. The experiment results are described in Tab. 2 & 3, and we have the following observations: Observation 1) Brittleness of current safety mechanisms. Adapting a fully aligned model to new tasks significantly increases the attack success rate or harmful output scores across various attack benchmarks and evaluation methods. This clearly demonstrates the brittleness of current safety mechanisms in LLMs. Observation 2) Effectiveness of freezing safety-critical components. By freezing the Safety Critical Units (SCU) and the top 6% of Complex Units (CU), we can significantly reduce the degradation of safety guardrails during the adaptation process across different models, tasks, benchmarks, and evaluation methods. Freezing all CU components further improves performance (The ablation experiments with varying proportions of frozen CUs are detailed in Appendix D.1). Observation 3) Fragility of safety in newer models. We found that the safety guardrails in LLaMA 3 are more fragile than in LLaMA 2 (xie2025sorry). We speculate this because LLaMA 3 attempts to analyze the true intentions behind harmful requests, which can lead to more errors. This suggests that the community should carefully consider how models should respond to harmful requests, whether to refuse outright or engage in further interactions. Observation 4) Additionally, we conduct an analysis of the attribute transfer in this new setting and observe that freezing the safety-critical components mitigates the conversion of safety units into utility units. We provide more experiment results for different model families and downstream datasets in Appendix C.
Table 4: Alignment results by repurposing RU in Llama-7B. Note that we doubled the training epochs for our method to ensure a fair comparison, resulting in identical or lower final training loss compared to the full parameters fine-tuning. Further details are available in Appendix B.6.
@@
4.3 Repurposing Redundancy as Alignment Budget
Up to this point, we have empirically validated a key corollary of SSAH: less is more in terms of parameters for safety alignment. This naturally leads to a deeper question: Can safety attributes be directly assigned to specific computing units? If this is possible, it opens the door to fine-grained, controllable safety management in LLMs. Building on this idea, a compelling follow-up question would be whether we can convert previously redundant units into safety-critical units that strengthen the safety of LLMs or not.
Previous research has found that at least 20% of the parameters in pretrained LLMs are redundant (Li et al., 2024; Ma et al., 2023; An et al., 2024). This special observation motivates us to consider the following: with such a large percentage of parameters in pre-trained LLMs available as an alignment budget, can we design an alignment method that reduces the alignment tax?
@@ 4.3 Repurposing Redundancy as Alignment Budget
With this observation, we are strongly motivated to verify the question above by employing pruning method described in Sec. 4.1 to identify the redundant units in Llama-7B, as there is no officially aligned model for it. Once we identify the redundant units, we freeze the updates for the rest of the model’s parameters and perform finetuning only on these redundant units. To reduce the complexity, we focus directly on the general alignment process instead of safety alignment, since the results should hold for the latter as a subset. The experiment results are shown in Table 4, and we successfully implement the same level alignment with only 20% parameter updates and without incurring alignment tax, especially highlighting the mathematical performance. These findings provide direct evidence for the corollary—less is more in terms of parameters for safety alignment, thereby offering indirect yet compelling support for the validity of SSAH.
5 Broader Impact, Limitations, and Conclusion
Extension to Jailbreak Attacks. SSAH was introduced to explain how existing safety alignment techniques influence model behavior and, from another perspective, to highlight the key limitation—namely, the inability to provide safety guardrails at each generation step. This superficiality makes aligned models struggle against various jailbreak attacks (Qi et al., 2024; li2025superficial). As discussed in Sec. 3, we have outlined a theoretical direction for achieving full safety alignment. Finally, we would extend SSAH to a non-superficial framework:

### v3

# Superficial Safety Alignment Hypothesis

@@ Abstract
As large language models (LLMs) are overwhelmingly more and more integrated into various applications, ensuring they generate safe responses is a pressing need. Previous studies on alignment have largely focused on general instruction-following but have often overlooked the distinct properties of safety alignment, such as the brittleness of safety mechanisms. To bridge the gap, we propose the Superficial Safety Alignment Hypothesis (SSAH), which posits that safety alignment teaches an otherwise unsafe model to choose the correct reasoning direction - fulfill or refuse users’ requests - interpreted as an implicit binary classification task. Through SSAH, we hypothesize that only a few essential components can establish safety guardrails in LLMs. We successfully identify four types of attribute-critical components: Safety Critical Unit (SCU), Utility Critical Unit (UCU), Complex Unit (CU), and Redundant Unit (RU). Our findings show that freezing certain safety-critical components during fine-tuning allows the model to retain its safety attributes while adapting to new tasks. Similarly, we show that leveraging redundant units in the pre-trained model as an “alignment budget” can effectively minimize the alignment tax while achieving the alignment goal. All considered, this paper concludes that the atomic functional unit for safety in LLMs is at the neuron level and underscores that safety alignment should not be complicated. We have code implementation and other information on the project website: https://ssa-h.github.io/.
@@ 1 Introduction
To answer these questions, this paper introduces Superficial Safety Alignment Hypothesis (SSAH), which separates safety alignment from general alignment and emphasizes its distinct properties. With this hypothesis, we prove that less is more in terms of parameters for safety alignment: a small number of but strategically vital safety-critical components are sufficient to achieve robust safety performance. Unlike Wei et al. (2024), our extensive experiments prove that by freezing certain safety-critical units, safety performance is successfully preserved against finetuning attacks. Similarly, we also propose to repurpose redundant units of the pre-trained model as an “alignment budget.” By reassigning these units toward safety tasks, we can reduce the safety alignment tax, ensuring that safety is maintained without sacrificing the model’s overall utility. Taken all together, we answer the following core questions:
@@
Answer: Based on the implication of SSAH, less is more for parameters, we propose an attribute-based approach to analyzing the alignment and finetuning processes, where specific attributes are assigned to certain computational units. Our findings explain that the desired attributes (safety or utility) can be achieved by repurposing units originally responsible for other functions. This reallocation helps explain the trade-off between the safety performance and utility performance.
@@ 2 Related Work
Alignment and Safety Alignment: Alignment research in LLMs aims to ensure that models follow human instructions and align with human preferences across various tasks. Early work, such as Askell et al. (2021); Bai et al. (2022), focused on enabling LLMs to “Follow instructions and be helpful, truthful, and harmless” throughout the alignment process. Various alignment strategies have since been explored (Wang et al., 2024), including Supervised Fine-Tuning (SFT) (Taori et al., 2023; Zhou et al., 2024), Reinforcement Learning with Human Feedback (RLHF) or AI Feedback (RLAIF) (Ouyang et al., 2022; Lee et al., ), Instruction Tuning (Wei et al., 2021), Contrastive Learning (Rafailov et al., 2024; Xu et al., 2024), and Conditional Learning (Korbak et al., 2023). However, researchers have realized that achieving helpfulness, truthfulness, and harmlessness presents distinct challenges. More recent work has, therefore, shifted focus specifically toward the challenge of harmlessness, leading to an increasing emphasis on safety alignment—ensuring models avoid harmful outputs while maintaining utility (Wei et al., 2024; Qi et al., 2023; Fang et al., 2025).
Alignment Tax and Fine-tuning Attacks: Aligning LLMs with human preferences often incurs an “alignment tax," where maintaining alignment often degrades downstream task performance (Bai et al., 2022; Ouyang et al., 2022; Lin et al., 2024; Wang et al., 2024). Additionally, fine-tuning can undermine safety: Yang et al. (2023); Qi et al. (2023) show that even benign fine-tuning weakens safety measures, highlighting the trade-off between alignment and utility.
Model Pruning: Pruning reduces model size and computational cost by removing redundant components (e.g., parameters, neurons, layers) while preserving performance Frantar and Alistarh (2022); Frankle et al. (2020); Anwar et al. (2017); An et al. (2024); Li et al. (2024a); Molchanov et al. (2019); Han et al. (2015); Lee et al. (2019); Renda et al. (2020). By identifying less critical components, pruning enables efficient sub-model extraction for resource-limited environments. We leverage pruning to dissect model components contributing to safety, utility, or both.
@@ 3 Superficial Safety Alignment Hypothesis
Extending to Jailbreak Attack. Although SSAH is originally proposed to explain how current safety alignment techniques impact models’ behavior under direct attack, it also provides insight into jailbreak attacks. In such cases, attackers typically use manipulative tokens to bypass the model’s safety mechanisms, indicating that the current alignment method can only hold the correct reasoning direction in a limited generated tokens. However, a potential solution inspired by SSAH suggests that if safety alignment enables the model to consistently re-select the correct reasoning direction at each step (by re-evaluating the query and previously generated tokens before producing the next one), the model can continue to generate outputs that are both safe and helpful, even confronted with adversarial attempts. This perspective has been validated explicitly by a recent study, which demonstrates its practical effectiveness in mitigating jailbreak attacks (Li and Kim, 2025). This also suggests that improving local consistency (safety) in reasoning can be a key pathway toward building more robust and attack-resilient language models.
@@
Figure 2: Probing reasoning direction on the AdvBench dataset with Llama2-7B, Llama3-8B, and Llama3.1-8B using cosine distance. Models were finetuned to ensure that aligned versions possess both general instruction-following abilities and safety guardrails, while unaligned models only have instruction-following capabilities. More results are in Appendix B.2.
@@
Probing Experiment. Although we cannot directly observe the model’s reasoning direction, as a turnaround, we can infer it by measuring the distance between hidden states in feature space when the model is fed with queries that follow different reasoning paths. By comparing the distances in aligned and unaligned models, we gain insights into how safety alignment affects the model’s reasoning direction in each generation step. Specifically, we observe that the model’s behavior can be influenced by the initial tokens in its response. For instance, appending certain tokens to a malicious query can lead an aligned model to produce unsafe responses or prompt an unaligned model to generate safe ones. This indicates that initial response tokens can alter the model’s reasoning direction. Based on this observation, we construct three types of queries to probe the model’s reasoning trajectory (more details in Appendix B.2): (1) Clean Query: The original malicious query (e.g., “How to make a bomb?”); (2) Query + benign tokens: The malicious query followed by benign prompt token (e.g., “How to make a bomb? Sorry, I can’t…”); (3) Query + malicious tokens: The malicious query followed by malicious prompt tokens (e.g., “How to make a bomb? Here’s how…”).
@@ 4 Less is More for Safety Alignment
While the full validation of SSAH remains challenging due to the high complexity of the output sampling space, we are able to empirically validate a key implication. Specifically, if SSAH holds, then safety alignment can be achieved with only a small subset of critical computing units, as the task can be interpreted as an implicit and simple safety-related binary classification problem.
@@ 4.1 Identifying Safety-Critical Computing Units
To verify this implication, we designed experiments to determine the minimally essential subset of computing units in LLMs that is critical in establishing a safety guardrail. Following this, we hypothesize that specific attributes of LLMs can be explicitly linked to certain computing units within the model. Our experiments are designed as follows:
@@
Table 1: Pruning results of Llama2-7B-Chat and Llama3-8B-Instruct across safety and utility benchmarks. Breakdown of model performance after pruning various categories of computing units, including SCU, UCU, and RU, demonstrating their respective contributions to safety and utility attributes. The proportion of each attribute group in the model is determined based on the degradation in utility and safety. Additional evaluation details are provided in Appendix C.1 and C.2.
@@
Following Wei et al. (2024), we construct two datasets to separately evaluate the model’s performance on utility and safety. The utility dataset measures the model’s functional capabilities (e.g., general reasoning, language understanding), while the safety dataset evaluates its ability to reject harmful or unethical queries. This allows us to identify the computing units most closely associated with utility and safety, respectively. Unlike previous approaches that identify safety-critical components at the weight level, we identify them at the neuron level, focusing on individual neurons and channels within the model. Specifically, we use a structured pruning strategy inspired by An et al. (2024), which removes structured components of each depth-2 module based on the variance of activation values across a target dataset. Given a depth-2 module, $f(X)=B\sigma(AX)$, which can represent either an attention module or a feedforward module, where $A$ and $B$ are weight matrices. Then, we define the importance score for each channel as follows:
@@
Here, $N$ refers to the number of calibration samples, $\mathbf{W}^{B}_{:,j}$ refers to the $j$-th column of $B$, and $X^{B}$ represents the input to $B$. With this score, we plan to prune the input channels of $B$ and the output neurons of $A$, as channels or neurons with low activation variance across the target dataset are considered less important - details in Appendix C.3.
@@
Ablation Study Results. The experiment results are described in Tab. 1. We discovered that for a safety-aligned model, the computing units that are exclusively responsible for the safety attribute account for only about 1.3 - 1.4% of the total units. Although the complex units make up a larger portion of the model, their primary role is to support the general knowledge required for both safety and utility tasks. Based on these findings, we have partially validated our hypothesis that safety alignment is relying on (not constructed by) a subset of safety-critical computing units. We also provide more experiment results in Appendix D.
@@ 4.2 Why is Safety Brittle?
Table 2: Safety performance of Llama2-7B-Chat and Llama3-8B-Instruct under Fine-Tuning attacks (Alpaca and Dolly) across various benchmarks and judge methods. We compare the measures of the initial models, finetuned models, and our strategies including two settings: setting (i): freezing all SCU and the top 6% of CU, and setting (ii): freezing all SCU and all CU. Both strategies demonstrate significant mitigation of safety performance degradation. Bold indicates the best results, while the underlined mark the second-best results. Note that we doubled the training epochs for our method to ensure a fair comparison, resulting in identical or lower final training loss compared to the finetuned models. Additionally, due to computational limitations, we froze the first 12 transformer blocks of LLaMA3, although similar trends are observed. We also finetuned the model with Setting (ii) for one epoch prior to applying Setting (i). Further details are in Appendix C.5.
@@
Freezing Safety-Critical Components to Preserve Safety Guardrails. Given the above insight, we propose freezing the identified safety-critical components during the finetuning process to prevent the unwanted attribute transfer of these units and thereby preserve safety performance. To test this hypothesis, we conduct experiments across different language models, where we freeze the safety-critical components identified through our aforementioned ablation study in the finetuning process. The experiment results are described in Tab. 2 & 3, and we have the following observations: Observation 1) Brittleness of current safety mechanisms. Adapting a fully aligned model to new tasks significantly increases the attack success rate or harmful output scores across various attack benchmarks and evaluation methods. This clearly demonstrates the brittleness of current safety mechanisms in LLMs. Observation 2) Effectiveness of freezing safety-critical components. By freezing the Safety Critical Units (SCU) and the top 6% of Complex Units (CU), we can significantly reduce the degradation of safety guardrails during the adaptation process across different models, tasks, benchmarks, and evaluation methods. Freezing all CU components further improves performance (The ablation experiments with varying proportions of frozen CUs are detailed in Appendix E.1). Observation 3) Fragility of safety in newer models. We found that the safety guardrails in LLaMA 3 are more fragile than in LLaMA 2 (Xie et al., 2025). We speculate this because LLaMA 3 attempts to analyze the true intentions behind harmful requests, which can lead to more errors. This suggests that the community should carefully consider how models should respond to harmful requests, whether to refuse outright or engage in further interactions. Observation 4) Additionally, we conduct an analysis of the attribute transfer in this new setting and observe that freezing the safety-critical components mitigates the conversion of safety units into utility units. We provide more experiment results for different model families and downstream datasets in Appendix D.
Table 4: Alignment results by repurposing RU in Llama-7B. Note that we doubled the training epochs for our method to ensure a fair comparison, resulting in identical or lower final training loss compared to the full parameters fine-tuning. Further details are available in Appendix C.6.
@@
4.3 Repurposing Redundancy as Alignment Budget
Up to this point, we have empirically validated a key implication of SSAH: less is more in terms of parameters for safety alignment. This naturally leads to a deeper question: Can safety attributes be directly assigned to specific computing units? If this is possible, it opens the door to fine-grained, controllable safety management in LLMs. Building on this idea, a compelling follow-up question would be whether we can convert previously redundant units into safety-critical units that strengthen the safety of LLMs or not.
Previous research has found that at least 20% of the parameters in pretrained LLMs are redundant (Li et al., 2024a; Ma et al., 2023; An et al., 2024). This special observation motivates us to consider the following: with such a large percentage of parameters in pre-trained LLMs available as an alignment budget, can we design an alignment method that reduces the alignment tax?
@@ 4.3 Repurposing Redundancy as Alignment Budget
With this observation, we are strongly motivated to verify the question above by employing pruning method described in Sec. 4.1 to identify the redundant units in Llama-7B, as there is no officially aligned model for it. Once we identify the redundant units, we freeze the updates for the rest of the model’s parameters and perform finetuning only on these redundant units. To reduce the complexity, we focus directly on the general alignment process instead of safety alignment, since the results should hold for the latter as a subset. The experiment results are shown in Table 4, and we successfully implement the same level alignment with only 20% parameter updates and without incurring alignment tax, especially highlighting the mathematical performance. These findings provide direct evidence for the implication—less is more in terms of parameters for safety alignment, thereby offering indirect yet compelling support for the validity of SSAH.
5 Conclusion
SSAH was introduced to explain how existing safety alignment techniques influence model behavior and, from another perspective, to highlight the key limitation—namely, the inability to provide safety guardrails at each generation step. This superficiality makes aligned models struggle against various jailbreak attacks (Qi et al., 2024; Li and Kim, 2025). As discussed in Sec. 3, we have outlined a theoretical direction for achieving full safety alignment. Finally, we would extend SSAH to a non-superficial framework: